# Kernel-based Koopman approximants for control: Flexible sampling, error analysis, and stability

### v1

@@ 3.2. Proportional bounds and asymptotic stability
Let $x^{*}\in\mathcal{X}$ be an equilibrium w.r.t. the dynamics Equation DS given by $F\in\mathcal{C}_{b}^{\lceil\sigma_{n,k}\rceil}(\Omega;\mathbb{R}^{n})$ with $k\geq 1$ and, thus, also of the data-driven surrogate Equation 10 represented by $\widehat{F}$. [^6] See Proposition 3.1. Here, we assume that $\psi_{i}(x)=x_{i}$ for $i\in[1:n]$. Furthermore, assume that one of the following conditions on the function $V:\Omega\to\mathbb{R}_{\geq 0}$ hold:
@@
    $V$ has a modulus of continuity $\omega_{V}\in\mathscr{K}$ satisfying the compatibility condition [^7] In case $\alpha_{V}$ is not some $m$-times continuously differentiable with $\alpha_{V}^{(m)}(0)\neq 0$, one may to replace Equation 20 with the more general condition $\exists\varepsilon>0:\limsup_{r\searrow 0}\ \frac{\omega_{V}(\varepsilon r)}{\alpha_{V}(r)}<1$.
@@
    $V(x)=\|x-x^{*}\|^{p}$ for some $p\in\mathbb{N}$, and (20) with $\omega_{V}(r)=r^{p}$.
@@
    Let $V$ be a Lyapunov function w.r.t. the dynamics Equation DS on $\Omega$ such that the decrease condition Equation 15 holds for all $x\in V^{-1}(c_{\Omega})$, which renders $V^{-1}(c_{\Omega})$ forward-invariant. Then, the equilibrium $x^{*}$ is asymptotically stable w.r.t. the data-driven surrogate Equation 10 for sufficiently-small fill distance $h_{\mathcal{X}}$.
@@
    The statement of assertion (i) holds upon switching the roles of $F$ and $\widehat{F}$, i.e., the existence of a Lyapunov function w.r.t. the data-driven surrogate dynamics Equation 10 implies asymptotic stability w.r.t. the dynamics Equation DS meaning that asymptotic stability is preserved.
@@
We begin with assertion (i). Hence, let $V$ be a Lyapunov function w.r.t. the dynamics Equation DS satisfying (a). We proceed along the lines of the proof of Theorem 3.5. Using the compatibility condition Equation 20 and the proportional bound of Corollary 3.8 in the estimate (16), we arrive at a counterpart of inequality Equation 17 with proportional left-hand side, i.e.,
$\omega_{V}(\bar{C}h_{\mathcal{X}}^{k-1/2}\|x-x^{*}\|)\leq s\alpha_{V}(\|x-x^{*}\|)$
holds for all $x\in V^{-1}(c_{\Omega})$ for sufficiently small fill distance. This allows to infer asymptotic stability of the equilibrium $x^{*}$ w.r.t. the data-driven surrogate dynamics Equation 10 using Proposition 3.4.
Let us now assume that (b) is satisfied. Then,
$V(\widehat{F}(x))$ $=V(\widehat{F}(x)-F(x)+F(x))=\|\widehat{F}(x)-F(x)+F(x)-x^{*}\|^{p}$
$\leq(\|\widehat{F}(x)-F(x)\|+\|F(x)-x^{*}\|)^{p}$
$=\sum_{j=0}^{p}\binom{p}{j}\|\widehat{F}(x)-F(x)\|^{p-j}\|F(x)-x^{*}\|^{j}.$
$\leq\|F(x)-x^{*}\|^{p}+\sum_{j=0}^{p-1}\binom{p}{j}\big{(}Ch_{\mathcal{X}}^{k-1/2}\|x-x^{*}\|\big{)}^{p-j}\max_{\xi\in\Omega}\|F^{\prime}(\xi)\|^{j}\|x-x^{*}\|^{j}$
$=V(F(x))+\sum_{j=0}^{p-1}\binom{p}{j}C^{p-j}h_{\mathcal{X}}^{(k-1/2)(p-j)}\max_{\xi\in\Omega}\|F^{\prime}(\xi)\|^{j}\|x-x^{*}\|^{p}$
$\leq V(x)-\alpha_{V}(\|x-x^{*}\|)+h_{\mathcal{X}}^{k-1/2}\Big{(}C+\max_{\xi\in\Omega}\|F^{\prime}(\xi)\|\Big{)}^{p}\|x-x^{*}\|^{p}$
$\leq V(x)-(1-s)\alpha_{V}(\|x-x^{*}\|)+h_{\mathcal{X}}^{k-1/2}\bar{C}^{p}\|x-x^{*}\|^{p}-s\alpha_{V}(\|x-x^{*}\|),$
where $\bar{C}=C+\max_{\xi\in\Omega}\|F^{\prime}(\xi)\|$ and $s\in(0,1)$. Thus, due to Equation 20 with $\omega_{V}(r)=r^{p}$, the assertion holds for sufficiently small fill distances $h_{\mathcal{X}}$. ∎
@@ 3.3. Example and numerical simulations
In this subsection, we illustrate our findings by means of an instructive example. Consider the nonlinear discrete-time system from [21] given by
@@
with state $x=(x_{1},x_{2})^{\top}$ in the compact set $\Omega=[-2,2]^{2}$. We show that $V(x)=\|x\|^{2}$ is a quadratic Lyapunov function on $\Omega$ for Equation 21 w.r.t. the equilibrium $x^{*}=0$, see also [21]. As we only have to verify the Lyapunov decrease condition Equation 15, we first compute
$V(F(x))$ $=$ $\frac{1}{64}\Big{[}1+(\|x\|^{2}-1)^{2}\Big{]}\|x\|^{2}=V(x)+\frac{1}{64}\Big{[}-63+(\|x\|^{2}-1)^{2}\Big{]}\|x\|^{2}.$
Since $\|x\|^{2}\leq 8$ holds for all $x\in\Omega=[-2,2]^{2}$, we have $(\|x\|^{2}-1)^{2}\leq 49$ on $\Omega$ and, thus, the desired Lyapunov decrease condition Equation 15 given by
$V(F(x))\leq V(x)-\frac{7}{32}\|x\|^{2}=V(x)-\alpha_{V}(\|x-x^{*}\|)$ (22)
with $\alpha_{V}(r)=7r^{2}/32$ on $\Omega$.
@@
We use the coordinate maps $\psi_{i}(x)=x_{i}$, $i\in\{1,2\}$, as observables in Equation 10 and set the smoothness degree of the Wendland kernels to $k=1$. As the Lyapunov function is given by the Euclidean norm, we can directly apply Theorem 3.9 using assumption (b) to conclude asymptotic stability of the data-driven surrogate dynamics Equation 12 and vice versa, nicely illustrating the suitability of the proposed data-driven surrogate dynamics Equation 12 for stability-related tasks. Correspondingly, when approximating $F$ by means of a regularized surrogate, that is Equation 10 with $\lambda>0$, we may apply Theorem 3.5 (and the comments thereafter) to deduce practical asymptotic stability.
@@
To validate our findings, we conduct numerical computations using a uniform grid $\mathcal{X}=\delta\mathbb{Z}^{2}\cap\Omega$ with mesh size $\delta>0$ leading to a fill distance $h_{\mathcal{X}}=2^{-1/2}\delta$. In addition, we consider a grid consisting of the Cartesian product of one-dimensional Chebyshev nodes.
@@
First, in Figure 1, we inspect the one-step prediction error of the surrogate model $\widehat{F}$ without regularization term, i.e. $\lambda=0$. The (uniform) validation grid $\mathcal{X}_{\text{val}}$ uses the mesh size $\delta_{\text{val}}=0.025$ and is given by $\mathcal{X}_{\text{val}}=\left(\delta_{\text{val}}\mathbb{Z}+\frac{\delta_{\text{val}}}{2}\right)^{2}\cap\Omega$. In the top row of Figure 1, the results for the uniform grid with mesh size $\delta\in\{0.2\cdot 2^{-i}\,|\,i\in[0:2]\}$ (and, therefore, $h_{\mathcal{X}}\in\{0.2\cdot 2^{-i-\frac{1}{2}}\,|\,i\in[0:2]\}$) can be seen. The intensity plots of the error $\|\widehat{F}(x)-F(x)\|_{2}$ for each point $x$ in the validation grid $\mathcal{X}_{\text{val}}$ show that the error decreases for decreasing fill distance. The lower row of Figure 1 depicts the results for the associated Chebyshev-based mesh with the same number of grid points $d\in\{441,1681,6561\}$ as used in the uniform grids. Again, we observe that the error decreases the more data points are chosen (as, correspondingly, the fill distance is decreased). In particular, the choice of a Chebyshev-based grid alleviates large errors at the boundary, as usual in interpolation.
@@
Figure 1. Approximation error $\|\widehat{F}(x)-F(x)\|_{2}$ for system Equation 21 with $d\in\{441,1681,6561\}$ data points for a uniform (mesh sizes $\delta\in\{0.2,0.1,0.05\}$; top) and a Chebychev-based grid (bottom).
@@
These also effect the maximum errors $\max\limits_{x\in\mathcal{X}_{\text{val}}}\|\widehat{F}(x)-F(x)\|_{2}$ that are depicted for the different numbers of data points $d$ and validation areas in Table 1. The closer we get to the origin of the domain, the smaller the maximal error, which indicates the proportional decrease proven in Corollary 2.8. It can also be seen that for smaller validation areas, i.e. the further afar the validation area is from the boundary, the better the surrogate model $\widehat{F}$ on the uniform grid performs in comparison to the Chebyshev-based grid. This is due to the construction of the Chebyshev-based grid that has a higher density of data points towards the boundary but therefore less data points close to the origin.
$d=441$ $d=1681$ $d=6561$
uniform Chebyshev uniform Chebyshev uniform Chebyshev
$[-2.0,2.0]^{2}$ 0.1205 0.1423 0.03770 0.00690 0.009540 0.000350
$[-1.0,1.0]^{2}$ 0.0053 0.0222 0.00030 0.00111 0.000021 0.000064
$[-0.5,0.5]^{2}$ 0.0007 0.0022 0.00004 0.00073 0.000001 0.000020
@@
Table 1. Maximal error for one-step prediction of System Equation 21 using the kEDMD surrogate $\widehat{F}$ for uniform and Chebyshev-based grid on different snippets of the validation grid.
@@
Next, we inspect regularized kEDMD with regularization parameter $\lambda=0.01$. In Figure 2, we again depict the one-step prediction errors $\|\widehat{F}_{\lambda}(x)-F(x)\|_{2}$ on $\Omega$ for the Chebyshev-based grid. The overall structure of the error is similar to what is depicted in Figure 1 without regularization. However, the closer we get to the origin, the error is increased by about two orders of magnitude compared to Figure 1. This is to be expected due to the additional term $\sqrt{\lambda}$ in the error estimate in Corollary 2.8.
Figure 2. Approximation error $\|\widehat{F}_{\lambda}(x)-F(x)\|_{2}$ for system Equation 21 with $d\in\{441,1681,6561\}$ data points for a Chebychev-based grid and regularization parameter $\lambda=0.01$.
@@
For a long-term evaluation of the errors, we choose a uniform grid $\mathcal{X}=\delta\mathbb{Z}^{2}\cap\Omega$ with $\delta=0.2\cdot 2^{-3}=0.025$ to learn the surrogate dynamics $\widehat{F}_{\lambda}$ for the regularization parameters $\lambda=0$ and $\lambda=0.01$. For the two initial conditions $x^{0}_{1}=r(1,0)^{\top}$ and $x^{0}_{2}=r(\cos(\pi/4),\sin(\pi/4))^{\top}$ that are located on the circle centered at the origin with radius $r=1.9$ the one-step errors $\|\widehat{F}_{\lambda}(x(k;x^{0}_{j}))-F(x(k;x^{0}_{j}))\|_{2}$ for $k\in[0:20]$ and $j\in\{1,2\}$ are visualized in Figure 3 in the image on the left. The difference of the one-step errors decays along the asymptotically stable trajectory of the original dynamics, confirming the proportional error bound Corollary 3.8 and the asymptotic stability of the surrogate model Theorem 3.9.
Lastly, we validate the Lyapunov decrease condition Equation 15, i.e., $V(\widehat{F}(x))-V(x)+\alpha_{V}(\|x\|)\leq 0$, where the data-driven surrogate dynamics $\widehat{F}$ of Equation 21 are generated using the mesh size $\delta=0.05$. In the right plot of Figure 3, the value of $-(V(\widehat{F}(x))-V(x)+\alpha_{V}(\|x\|))$ for $x\in\mathcal{X}_{\text{val}}$ is plotted and we can observe that the decrease condition is in fact preserved as stated in Theorem 3.9.
@@
Figure 3. Left: One-step error $\|\widehat{F}_{\lambda}(x(k;x^{0}_{j}))-F(x(k;x^{0}_{j}))\|_{2}$ along the trajectory of Equation 21 for $r=1.9$ and $\tilde{r}=0.95\sqrt{2}$ with different regularization parameters $\lambda$. Right: Visualization of the Lyapunov decrease $-(V(\widehat{F}(x))-V(x)+\alpha_{V}(\|x\|))$ for system Equation 21 with the kEDMD surrogate $\widehat{F}$.

### v3

@@ 3.2. Proportional bounds and asymptotic stability
Let $x^{*}\in\mathcal{X}$ be an equilibrium w.r.t. the dynamics eq. DS given by $F\in\mathcal{C}_{b}^{\lceil\sigma_{n,k}\rceil}(\Omega;\mathbb{R}^{n})$ with $\sigma_{n,k}$ defined by eq. 3 and $k\geq 1$ and, thus, also of the data-driven surrogate eq. 11 represented by $\widehat{F}$. [^6] See proposition 3.1. Furthermore, let $\psi_{i}(x)=x_{i}$, $i\in[1:n]$, be given. In addition, suppose that one of the following conditions on the function $V:\Omega\to\mathbb{R}_{\geq 0}$ hold:
@@
[…]
@@
    $V(x)=\|x-x^{*}\|^{p}$ for some $p\in\mathbb{N}$, and $\alpha_{V}(r)=c_{V}r^{p}$. [^8] Clearly, this implies also the compatibility condition eq. 20.
@@
    Let $V$ be a Lyapunov function w.r.t. the dynamics eq. DS on $\Omega$ such that the decrease condition eq. 15 holds for all $x\in V^{-1}(c_{\Omega})$, which renders $V^{-1}(c_{\Omega})$ forward-invariant. Then, the equilibrium $x^{*}$ is asymptotically stable w.r.t. the data-driven surrogate eq. 9 for sufficiently-small fill distance $h_{\mathcal{X}}$.
@@
    The statement of assertion (i) holds upon switching the roles of $F$ and $\widehat{F}$, if the comparison functions of the Lyapunov function are uniform for all $h_{\mathcal{X}}\in(0,h_{0}]$ for some $h_{0}$, i.e., the existence of a Lyapunov function w.r.t. the data-driven surrogate dynamics eq. 9 uniform in the fill distance implies asymptotic stability w.r.t. the dynamics eq. DS meaning that asymptotic stability is preserved.
@@
We begin with assertion (i). Let $V$ be a Lyapunov function w.r.t. eq. DS satisfying condition (a). We proceed along the lines of the proof of theorem 3.5. Using the compatibility condition eq. 20 and the proportional bound of corollary 3.8 in the estimate (17), we arrive at a counterpart of inequality eq. 18 with proportional left-hand side, i.e., $\omega_{V}(\bar{C}h_{\mathcal{X}}^{k-1/2}\|x-x^{*}\|)\leq s\alpha_{V}(\|x-x^{*}\|)$ holds for all $x\in V^{-1}(c_{\Omega})$ for sufficiently small fill distance. This allows to infer asymptotic stability of the equilibrium $x^{*}$ w.r.t. the data-driven surrogate dynamics eq. 9 using proposition 3.4.
Let us now assume that (b) holds. Then, using $J_{F}$ to denote the Jacobian of $F$,
$V(\widehat{F}(x))$ $=$ $\|\widehat{F}(x)-F(x)+F(x)-x^{*}\|^{p}\leq\Big(\|\widehat{F}(x)-F(x)\|+\|F(x)-x^{*}\|\Big)^{p}$
$=$ $\sum\nolimits_{j=0}^{p}\binom{p}{j}\|\widehat{F}(x)-F(x)\|^{p-j}\|F(x)-x^{*}\|^{j}.$
$\leq$ $\|F(x)-x^{*}\|^{p}+\sum_{j=0}^{p-1}\binom{p}{j}\left(Ch_{\mathcal{X}}^{k-1/2}\|x-x^{*}\|\right)^{p-j}\max_{\xi\in\Omega}\|J_{F}(\xi)\|^{j}\|x-x^{*}\|^{j}$
$\leq$ ${V(F(x))+h_{\mathcal{X}}^{k-1/2}\hskip-2.84526pt\underbrace{\left[\sum_{j=0}^{p-1}\binom{p}{j}C^{p-j}h_{0}^{(k-\frac{1}{2})(p-j-1)}\max_{\xi\in\Omega}\|J_{F}(\xi)\|^{j}\right]}_{=:\bar{C}}\hskip-2.84526pt\|x-x^{*}\|^{p}}$
[…]
$\leq$ $V(x)-{\Big(1-h_{\mathcal{X}}^{k-1/2}\bar{C}/c_{V})\Big)\cdot\alpha_{V}(\|x-x^{*}\|)},$
where $s\in(0,1)$. Thus, the assertion follows for sufficiently small fill distances $h_{\mathcal{X}}$.
The second statement (ii) follows, analously as in theorem 3.5, upon switching the roles of $\hat{F}$ and $F$ using uniformity of the Lyapunov function for small fill distances. ∎
@@ 3.3. Example and numerical simulations
We illustrate our findings by an instructive example. Consider the nonlinear discrete-time system from [29] given by
@@
with state $x=(x_{1},x_{2})^{\top}$ in the compact set $\Omega=[-2,2]^{2}$, for which $V(x)=\|x\|^{2}$ serves as quadratic Lyapunov function on $\Omega$ w.r.t. the equilibrium $x^{*}=0$, see [29].
@@
We use the coordinate maps $\psi_{i}(x)=x_{i}$, $i\in\{1,2\}$, as observables in eq. 9 and set the smoothness degree $k=1$ of the Wendland kernels. As the Lyapunov function is given by the Euclidean norm, we can directly apply theorem 3.9 using assumption (b) to conclude asymptotic stability of the data-driven surrogate dynamics eq. 11 and vice versa. When approximating $F$ by means of a regularized surrogate ($\lambda>0$), we apply theorem 3.5 (and the comments thereafter) to deduce practical asymptotic stability.
@@
To validate our findings, we conduct numerical computations using a uniform grid $\mathcal{X}=\delta\mathbb{Z}^{2}\cap\Omega$ with mesh size $\delta>0$ leading to a fill distance $h_{\mathcal{X}}=2^{-1/2}\delta$. In addition, we consider a grid consisting of Padua points as described in [7]. For degree $\gamma_{\text{P}}\in\mathbb{N}$, the Padua points of family $s=1$ are given by the set of points
$\text{Pad}_{\gamma_{P}}=\Big\{P\left(\frac{k\pi}{\gamma_{\text{P}}(\gamma_{\text{P}}+1)}\right)\mid k=0,\dots,\gamma_{\text{P}}(\gamma_{\text{P}}+1)\Big\},$
@@
where $P:[0,\pi]\rightarrow\mathbb{R}^{2}$ is defined by $P(t)=\left(-\cos((\gamma_{\text{P}}+1)t)\ -\cos(\gamma_{\text{P}}t)\right)^{\top}$. This results in $|\text{Pad}_{\gamma_{P}}|=(\gamma_{P}+1)(\gamma_{P}+2)/2$ grid points. Using a grid of Padua points reduces the condition number for two-dimensional problems by minimizing the Lebesgue constant, similar to Chebychev nodes in one dimension. First, in fig. 1, we inspect the one-step prediction error of the surrogate model $\widehat{F}$ without regularization term, i.e. $\lambda=0$. The (uniform) validation grid $\mathcal{X}_{\text{val}}$ uses the mesh size $\delta_{\text{val}}=0.025$ and is given by $\mathcal{X}_{\text{val}}=\left(\delta_{\text{val}}\mathbb{Z}+\frac{\delta_{\text{val}}}{2}\right)^{2}\cap\Omega$. In the top row of fig. 1, the results for the uniform grid with mesh size $\delta\in\{0.2\cdot 2^{-i}\,|\,i\in[0:2]\}$ (and, therefore, $h_{\mathcal{X}}\in\{0.2\cdot 2^{-i-\frac{1}{2}}\,|\,i\in[0:2]\}$) can be seen. The intensity plots of the error $\|\widehat{F}(x)-F(x)\|_{2}$ for each point $x$ in the validation grid $\mathcal{X}_{\text{val}}$ show that the error decreases for decreasing fill distance. The lower row of fig. 1 depicts the results for the associated Padua-based mesh with degrees $\gamma_{\text{P}}\in\{28,56,113\}$, which results in $d\in\{435,1653,6555\}$ grid points. Again, we observe that the error decreases the more data points are chosen (as, correspondingly, the fill distance is decreased). In particular, the choice of a Padua-based grid alleviates large errors at the boundary.
@@
Figure 1. Approximation error $\|\widehat{F}(x)-F(x)\|_{2}$ for system eq. 21 with $d\in\{441,1681,6561\}$ data points for a uniform (mesh sizes $\delta\in\{0.2,0.1,0.05\}$; top) and a Padua grid (middle), $d\in\{436,1654,6556\}$. Bottom: Error $\|\widehat{F}_{\lambda}(x)-F(x)\|_{2}$ with regularization parameter $\lambda=0.01$.
@@
This also effect the maximum errors $\max_{x\in\mathcal{X}_{\text{val}}}\|\widehat{F}(x)-F(x)\|_{2}$, see table 1. On the full domain, the usage of the Padua grid reduces the maximal error visibly. The closer we get to the origin of the domain, the smaller the maximal error, which indicates the proportional decrease proven in corollary 2.8. It can also be seen that for smaller validation areas, i.e. the further afar the validation area is from the boundary, the better the surrogate model $\widehat{F}$ on the uniform grid performs, and produces errors comparable to the ones obtained from the Padua-based grid.
$d=441$ $d=435$ $d=1681$ $d=1653$ $d=6561$ $d=6555$
uniform Padua uniform Padua uniform Padua
$2.0,2.0]^{2}$ 0.1205 0.0127 0.03770 0.00079 0.009540 0.0001500
$1.0,1.0]^{2}$ 0.0053 0.0044 0.00030 0.00033 0.000021 0.0000380
$5,0.5]^{2} 0.0$007 0.0008 0.00004 0.00002 0.000001 0.0000009
@@
Table 1. Maximal error for one-step prediction of System eq. 21 using the kEDMD surrogate $\widehat{F}$ for uniform and Padua-based grid on different snippets of the validation grid.
@@
Next, we inspect regularized kEDMD with regularization parameter $\lambda=0.01$. While the results resemble the ones depicted in fig. 1 without regularization, the error is increased by about two orders of magnitude close to the origin due to the additional term $\sqrt{\lambda}$ in corollary 2.8.
@@
For a long-term evaluation of the errors, we choose a uniform grid $\mathcal{X}=\delta\mathbb{Z}^{2}\cap\Omega$ with $\delta=0.2\cdot 2^{-3}=0.025$ to learn the surrogate dynamics $\widehat{F}_{\lambda}$ for the regularization parameters $\lambda=0$ and $\lambda=0.01$. For the two initial conditions $x^{0}_{1}=r(1,0)^{\top}$ and $x^{0}_{2}=r(\cos(\pi/4),\sin(\pi/4))^{\top}$ that are located on the circle centered at the origin with radius $r=1.9$ the one-step errors $\|\widehat{F}_{\lambda}(x(k;x^{0}_{j}))-F(x(k;x^{0}_{j}))\|_{2}$ for $k\in[0:20]$ and $j\in\{1,2\}$ are visualized in fig. 2 in the image on the left. The difference of the one-step errors decays along the asymptotically stable trajectory of the original dynamics, confirming the proportional error bound corollary 3.8 and the asymptotic stability of the surrogate model theorem 3.9. Lastly, we validate the Lyapunov decrease condition eq. 15, i.e., $V(\widehat{F}(x))-V(x)+\alpha_{V}(\|x\|)\leq 0$, where the data-driven surrogate dynamics $\widehat{F}$ of eq. 21 are generated using the mesh size $\delta=0.05$. In the right plot of fig. 2, the value of $(V(\widehat{F}(x))-V(x)+\alpha_{V}(\|x\|))$ for $x\in\mathcal{X}_{\text{val}}$ is plotted and we can observe that the decrease condition is in fact preserved as stated in theorem 3.9.
@@
Figure 2. Left: $\|\widehat{F}_{\lambda}(x(k;x^{0}_{j}))-F(x(k;x^{0}_{j}))\|_{2}$ along trajectories of eq. 21 for $r=1.9$, $\tilde{r}=0.95\sqrt{2}$. Right: Lyapunov decrease $V(x)-V(\widehat{F}(x))-\alpha_{V}(\|x\|)$ for system eq. 21 with the kEDMD surrogate $\widehat{F}$.